May you please help me with this problem, I am trying to find the value of "y" for the given value of "x" in each of the following equations. Then write the ordered pairs.
x+y=10
x-y=8

Answer by AlgebraGirl08(3)

Hi, if you're trying to find the value of "y" for the given value of "x", they should give you a value of "x". If you have that, then you just plug that number into the equation wherever "x" is.
Examples...
-1- x+y=10 x=5
You want to plug 5 in for the "x" value.
(5)+y=10
Then you subtract the 5 from both sides, from the five and the ten, to get the "y" alone.
((y=5))
And that's your answer.
-2- x-y=8 x=2
In this example you're dealing with negatives. Again, first you want to plug 2 into the equation.
(2)-y=8
Then to get the "y" alone you divide the 2 from both sides, the 2 and the 8.
-y=4
Now, there's one thing that is different from this problem and the one before it. The negative sign. The negative sign signifies that the value of "y", whatever it may be, isn't positive, it's negative. -y is also -1y, because you have 1 negative "y". What you wanna do is treat the negative sign just like a number and divide it from both sides.
-1y=4
((y=-4))